OpenStudy (anonymous):

Help please!!! A recipe that makes 7 servings calls for 1 and 1/6 cups of juice. How many cups of juice are there per serving? A. 1/6 B. 1/7 C. 7/6 D. 6/7

OpenStudy (aripotta):

so you wanna convert 1 and 1/6 into an improper fraction. once you do that you can divide it by 7

OpenStudy (aripotta):

and to divide fractions, you go by the saying "same, change, flip" meaning you keep the first fraction the same, you change the division to multiplication, and flip the second fraction so\[\frac{ w }{ x }\div \frac{ y }{ z } = \frac{ w }{ x }\times \frac{ z }{ y }\]

OpenStudy (anonymous):

So 7/6 *7?

OpenStudy (aripotta):

7 = 7/1 right? and we have to flip the second fraction. so it would be 7/6 x 1/7

OpenStudy (anonymous):

I'm pretty sure it's A) 1/6

OpenStudy (anonymous):

yes and that = 7/42 then you simplify

OpenStudy (anonymous):

Right?

OpenStudy (aripotta):

yep

OpenStudy (anonymous):

So 1/6

OpenStudy (aripotta):

correctomundo

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!!

OpenStudy (aripotta):

no problemo :)